OpenStudy (anonymous):

Helping son with homework and am stumped myself. The answer key has a different answer than I figured. Can someone please show me the work to solve the following problem? If the height of a rectangle is 2/5mm and the perimeter is 20 2/15mm, what is the length of the base of the rectangle? Thank you.

OpenStudy (anonymous):

p = 2(l+b) 20 2/15 = 2(2/5 + l)

OpenStudy (anonymous):

perimeter = 2 ( length + height) 20 2/15 = 2(2/5+length)

OpenStudy (anonymous):

l= 10 1/15 - 2/5

OpenStudy (anonymous):

145/15 = 29/3 = 9 2/3

OpenStudy (anonymous):

Thank you.